# Matrix Rosenthal and Concentration Inequalities for Markov Chains with Applications in Statistical Learning

### v1

@@ 4.1 Proof of the Rosenthal-Burkholder Inequality for Matrix Martingales
$\gamma\epsilon=2d\left({\frac{e}{N}\frac{\delta_{1}^{2}}{\delta_{2}^{2}}}\right)^{N}\beta^{p}<1,\quad\text{ where }\ N=\frac{\beta-(1+\delta_{2})}{\delta_{2}}.$ (10)
@@
$\left\|{\sup_{k\in{\mathbb{N}}}\|{\bm{M}}_{k}\|}\right\|_{p}\leq 5.01\!\left({1{+}e^{-p/c}{+}\frac{1}{c{\vee}\log d}}\right)\left({\sqrt{c{\vee}\log d}\ e^{p{/}c}\left\|{\left\|{\left\langle{{\bm{M}}}\right\rangle_{\infty}}\right\|^{1{/}2}}\right\|_{p}{+}(c{\vee}\log d)\left\|{\sup_{k\in{\mathbb{N}}}\left\|{{\bm{X}}_{k}}\right\|}\right\|_{p}}\right).$ (11)
@@ 4.2 Proof of the Matrix Rosenthal Inequality for Markov Chains
    Under Assumption 2.2 and $\|\|{\bm{F}}\|\|_{V^{\alpha}}<\infty$ for some $\alpha\in(0,1]$, we have that ${\bm{G}}=\sum_{k=0}^{\infty}{\mathrm{Q}}^{k}{\bm{F}}$ is a solution to the Poisson equation and satisfies $\|{\bm{G}}(z)\|\leq[\varkappa^{\alpha}2^{3-\alpha}/(3\alpha)](\pi(V)+V(z))^{\alpha}t_{\operatorname{mix}}\|\|{\bm{F}}\|\|_{V^{\alpha}}$ for any $z\in{\mathcal{Z}}$.
@@
where ${\bm{X}}_{k}={\bm{G}}(Z_{k})-({\mathrm{Q}}{\bm{G}})(Z_{k-1})$ is a $({\mathscr{F}}_{k})_{k\in{\mathbb{N}}}$-adapted martingale difference sequence, and we denote ${\bm{M}}_{k}=\sum_{j=1}^{k}{\bm{X}}_{j}$ as the related martingale and $\langle{\bm{M}}\rangle_{k}=\sum_{j=1}^{k}{\mathbb{E}}_{j-1}[({\bm{X}}_{j})^{2}]$ as the quadratic variation. We define ${\bm{M}}_{0}={\bm{X}}_{0}=\bm{0}_{d\times d}$.
As in the proof of the Hilbert space-valued Rosenthal inequality for ergodic Markov chains (Durmus et al., 2023), we first deal with stationary Markov chains, and then generalize the results to non-stationary chains.
@@
We assume $(Z_{k})_{k\in{\mathbb{N}}}$ is a stationary Markov chain, namely $Z_{0}\sim\pi$. By Minkowski inequality and Lemma 4.2, for all $p\geq 2$, we have
$\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\pi,p}\leq\left\|{\sup_{k\leq n}\left\|{{\bm{M}}_{k}}\right\|}\right\|_{\pi,p}+2\left\|{\left\|{{\bm{G}}(Z_{0})}\right\|}\right\|_{\pi,p}.$ (13)
@@
$\left\|{\sup_{k\leq n}\left\|{{\bm{M}}_{k}}\right\|}\right\|_{\pi,p}\leq$ $C_{R,1}\sqrt{p\vee\log d}\Big{\|}\left\|{\left\langle{{\bm{M}}}\right\rangle_{n}}\right\|^{1/2}\Big{\|}_{\pi,p}+C_{R,2}(p\vee\log d)\left\|{\sup_{k\leq n}\|{\bm{X}}_{k}\|}\right\|_{\pi,p},$ (14)
@@
Now, we aim to give an upper bound for $\|\|\langle{\bm{M}}\rangle_{n}\|^{1/2}\|_{\pi,p}$. Note that $\langle{{\bm{M}}}\rangle_{n}=\sum_{k=1}^{n}{\mathbb{E}}_{k-1}[({\bm{X}}_{k})^{2}]$ and
@@
where we defined ${\bm{H}}(z)=({\mathrm{Q}}{\bm{G}}^{2})(z)-[({\mathrm{Q}}{\bm{G}})(z)]^{2}$ for any $z\in{\mathcal{Z}}$. According to the Poisson equation Eqn. (12) and the definition of ${\bm{\Sigma}}_{\pi}({\bm{F}})$ (Eqn. (2)), we have
$\pi\left({{\bm{H}}}\right)=$ $\pi\left({{\mathrm{Q}}{\bm{G}}^{2}-({\mathrm{Q}}{\bm{G}})^{2}}\right)$
$=$ $\pi\left({{\bm{G}}^{2}-\left({{\bm{G}}-{\bm{F}}}\right)^{2}}\right)$
$=$ $\pi\left({{{\bm{F}}}{\bm{G}}+{\bm{G}}{\bm{F}}-{{\bm{F}}}^{2}}\right)$
$=$ $\pi\left({{{\bm{F}}}^{2}}\right)+\sum_{k=1}^{\infty}\pi\left({{{\bm{F}}}{\mathrm{Q}}^{k}{{\bm{F}}}+\left({{\mathrm{Q}}^{k}{{\bm{F}}}}\right){{\bm{F}}}}\right)$
$=$ ${\bm{\Sigma}}_{\pi}({\bm{F}}).$
@@
$\Big{\|}\left\|{\left\langle{{\bm{M}}}\right\rangle_{n}}\right\|^{1/2}\Big{\|}_{\pi,p}=$ $\sqrt{\left\|{\left\|{\sum_{k=1}^{n}{\bm{H}}(Z_{k-1})}\right\|}\right\|_{\pi,p/2}}$ (15)
$\leq$ $\sqrt{\left\|{n\left\|{\pi({\bm{H}})}\right\|+\left\|{\sum_{k=1}^{n}\left({{\bm{H}}(Z_{k-1})-\pi({\bm{H}})}\right)}\right\|}\right\|_{\pi,p/2}}$
$\leq$ $\sqrt{n\left\|{{\bm{\Sigma}}_{\pi}({\bm{F}})}\right\|+\left\|{\left\|{\sum_{k=0}^{n-1}\bar{\bm{H}}(Z_{k})}\right\|}\right\|_{\pi,p/2}}$
$\leq$ $\sqrt{n}\left\|{{\bm{\Sigma}}_{\pi}({\bm{F}})}\right\|^{1/2}+\left\|{\left\|{\sum_{k=0}^{n-1}\bar{\bm{H}}(Z_{k})}\right\|}\right\|_{\pi,p/2}^{1/2},$
where we denoted $\bar{{\bm{H}}}={{\bm{H}}}-\pi({\bm{H}})$.
@@
$\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\pi,p}\leq$ $C_{R,1}\sqrt{n(p\vee\log d)}\left\|{{\bm{\Sigma}}_{\pi}({\bm{F}})}\right\|^{1/2}+C_{R,1}\sqrt{p\vee\log d}\left\|{\left\|{\sum_{k=0}^{n-1}\bar{\bm{H}}(Z_{k})}\right\|}\right\|_{\pi,p/2}^{1/2}$ (16)
$+C_{R,2}(p\vee\log d)\left\|{\sup_{k\leq n}\|{\bm{X}}_{k}\|}\right\|_{\pi,p}+2\left\|{\left\|{{\bm{G}}(Z_{0})}\right\|}\right\|_{\pi,p},$
@@
If we assume Assumption 2.1 and ${\bm{F}}^{2}(z)\preccurlyeq{\bm{\Upsilon}}$, then by Lemma 4.2 (a), it holds that $\left({{\mathrm{Q}}{\bm{G}}^{2}}\right)(Z_{k-1})\preccurlyeq(64/9)t_{\operatorname{mix}}^{2}{\bm{\Upsilon}}$, thus $\langle{{\bm{M}}}\rangle_{n}\preccurlyeq(64/9)nt_{\operatorname{mix}}^{2}{\bm{\Upsilon}}$. By Theorem 2.1 (b), the inequality (14) still holds if we replace $d$ by $d_{n,p}=e\cdot r([(64/9)t_{\operatorname{mix}}^{2}{\bm{\Upsilon}}]\wedge[(1/n)\|\|\left\langle{{\bm{M}}}\right\rangle_{n}\|\|_{\pi,\frac{p}{2}}])$, and double the right-hand side. As $n\to\infty$, by ergodicity and the fact that $\pi({\bm{H}})={\bm{\Sigma}}_{\pi}({\bm{F}})$, we find that $d_{n,p}\to e\cdot r([(64/9)t_{\operatorname{mix}}^{2}{\bm{\Upsilon}}]\wedge\|{\bm{\Sigma}}_{\pi}({\bm{F}})\|)$, which is a natural dimension factor.
@@
    $\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\pi,p}\leq$ $\frac{16}{3}C_{R,1}\sqrt{(p\vee\log d)nt_{\operatorname{mix}}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}+\left({\frac{16}{3}C_{R,2}(p\vee\log d)+\frac{19}{3}}\right)t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty},$
    Furthermore, if there exists an ${\bm{\Upsilon}}\in{\mathbb{H}}^{d}_{+}$ such that ${\bm{F}}^{2}(z)\preccurlyeq{\bm{\Upsilon}}$ for all $z\in{\mathcal{Z}}$, then the above inequality also holds when we replace $d$ by $e\cdot r([(64/9){\bm{\Upsilon}}]\wedge[(1/\lfloor{n}/{t_{\operatorname{mix}}}\rfloor)\|\|\langle{{\bm{M}}^{(t_{\operatorname{mix}})}}\rangle_{\lfloor{n}/{t_{\operatorname{mix}}}\rfloor}\|\|_{\pi,\frac{p}{2}}])$, and double the right-hand side, where ${\bm{M}}^{(t_{\operatorname{mix}})}$ is a matrix martingale defined in Appendix B.2.
@@
    $\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\pi,p}\leq$ $\frac{8}{3}C_{R,1}p\sqrt{(p\vee\log d)nt_{\operatorname{mix}}}\left[{\varkappa\pi(V)}\right]^{1/p}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{1/p}}$
    $+\left({\frac{16}{3}C_{R,2}(p\vee\log d)+\frac{11}{3}}\right)pt_{\operatorname{mix}}n^{1/p}\left[{\varkappa\pi(V)}\right]^{1/p}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{1/p}}.$
@@
$C_{R,1}\sqrt{p\vee\log d}\left\|{\left\|{\sum_{k=0}^{n-1}\bar{\bm{H}}(Z_{k})}\right\|}\right\|_{\pi,p/2}^{1/2}\lesssim\left({p\vee\log d}\right)^{3/4}t_{\operatorname{mix}}^{5/4}n^{1/4}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}+\left({p\vee\log d}\right)t_{\operatorname{mix}}^{3/2}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}.$ (17)
And it is easy to see that the other two residual terms are dominated by the term. If we further assume ${\bm{F}}^{2}(z)\preccurlyeq{\bm{\Upsilon}}$, then by Lemma 4.2 (a), the above inequality still holds if we replace $d$ introduced in the l.h.s. by $e\cdot r([(64/9)t_{\operatorname{mix}}^{2}{\bm{\Upsilon}}]\wedge[(1/n)\|\|\left\langle{{\bm{M}}}\right\rangle_{n}\|\|_{\pi,\frac{p}{2}}])$, and $d$ introduced in the r.h.s. (see Lemma 4.3) by $e\cdot r([(64/9)t_{\operatorname{mix}}^{2}{\bm{\Upsilon}}]\wedge[(1/\lfloor{n}/{t_{\operatorname{mix}}}\rfloor)\|\|\langle{{\bm{M}}^{(t_{\operatorname{mix}})}_{{\bm{H}}}}\rangle_{\lfloor{n}/{t_{\operatorname{mix}}}\rfloor}\|\|_{\pi,\frac{p}{4}}])$, and double the r.h.s., where the matrix martingale ${\bm{M}}^{(t_{\operatorname{mix}})}_{\bar{\bm{H}}}$ is defined in Appendix B.2 with $\bm{F}$} replaced by $\bar{\bm{H}}$.
@@
$\leq$ $\sup_{z\in{\mathcal{Z}}}\frac{[p\varkappa^{1/p}2^{3-1/p}/3](\pi(V)+V(z))^{1/p}t_{\operatorname{mix}}\|\|{\bm{F}}\|\|_{V^{1/p}}}{V^{1/p}(z)}$
$\lesssim$ $p\left[{\varkappa\pi(V)}\right]^{1/p}t_{\operatorname{mix}}\|\|{\bm{F}}\|\|_{V^{1/p}},$
@@
$C_{R,1}\sqrt{p\vee\log d}\left\|{\left\|{\sum_{k=0}^{n-1}\bar{\bm{H}}(Z_{k})}\right\|}\right\|_{\pi,p/2}^{1/2}\lesssim$ $p^{3/2}\left({p\vee\log d}\right)^{3/4}t_{\operatorname{mix}}^{5/4}n^{1/4}\left[{\varkappa\pi(V)}\right]^{2/p}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{1/p}}$
$+p^{3/2}\left({p\vee\log d}\right)t_{\operatorname{mix}}^{3/2}n^{1/p}\left[{\varkappa\pi(V)}\right]^{2/p}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{1/p}}.$
@@
$\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\xi,p}\leq\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\pi,p}+\frac{2}{n}\left({1+128(t_{\operatorname{mix}}/\log 4)^{p}(p+1)^{p+1/2}e^{-p}}\right)^{1/p}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}.$
@@
$\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\xi,p}\leq\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\pi,p}+\frac{32}{15n}\varkappa^{1/p}\left({\xi(V)+\pi(V)}\right)^{1/p}pt_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{1/p}}.$

### v2

@@ 4.1 Proof of the Rosenthal-Burkholder Inequality for Matrix Martingales
[…]
@@
[…]
@@ 4.2 Proof of the Matrix Rosenthal Inequality for Markov Chains
    Under Assumption 2.2 and $\|\|{\bm{F}}\|\|_{V^{\alpha}}<\infty$ for some $\alpha\in(0,1]$, we have that ${\bm{G}}=\sum_{k=0}^{\infty}{\mathrm{Q}}^{k}{\bm{F}}$ is a solution to the Poisson equation and satisfies $\|{\bm{G}}(z)\|\leq[\varkappa^{\alpha}2^{3-\alpha}/(3\alpha)](\mu(V)+V(z))^{\alpha}t_{\operatorname{mix}}\|\|{\bm{F}}\|\|_{V^{\alpha}}$ for any $z\in{\mathcal{Z}}$.
@@
where ${\bm{X}}_{k}={\bm{G}}(Z_{k})-({\mathrm{Q}}{\bm{G}})(Z_{k-1})$ is a $({\mathscr{F}}_{k})_{k\in{\mathbb{N}}}$-adapted martingale difference sequence, and we denote ${\bm{M}}_{k}=\sum_{j=1}^{k}{\bm{X}}_{j}$ as the related martingale and $\langle{\bm{M}}\rangle_{k}=\sum_{j=1}^{k}{\mathbb{E}}_{j-1}[({\bm{X}}_{j})^{2}]$ as the quadratic variation. We define ${\bm{M}}_{0}={\bm{X}}_{0}=\bm{0}_{d\times d}$. We first deal with stationary Markov chains, and then generalize the results to non-stationary chains.
@@
We assume $(Z_{k})_{k\in{\mathbb{N}}}$ is a stationary Markov chain, namely $Z_{0}\sim\mu$. By Minkowski inequality and Lemma 4.2, for all $p\geq 2$, we have
[…]
@@
$\left\|{\sup_{k\leq n}\left\|{{\bm{M}}_{k}}\right\|}\right\|_{\mu,p}\leq$ $C_{1}\sqrt{p\vee\log d}\Big\|\left\|{\left\langle{{\bm{M}}}\right\rangle_{n}}\right\|^{1/2}\Big\|_{\mu,p}+C_{2}(p\vee\log d)\left\|{\sup_{k\leq n}\|{\bm{X}}_{k}\|}\right\|_{\mu,p},$ (16)
@@
Now, we aim to give an upper bound for $\|\|\langle{\bm{M}}\rangle_{n}\|^{1/2}\|_{\mu,p}$. Note that $\langle{{\bm{M}}}\rangle_{n}=\sum_{k=1}^{n}{\mathbb{E}}_{k-1}[({\bm{X}}_{k})^{2}]$ and
@@
where we defined ${\bm{H}}(z)=({\mathrm{Q}}{\bm{G}}^{2})(z)-[({\mathrm{Q}}{\bm{G}})(z)]^{2}$ for any $z\in{\mathcal{Z}}$. According to the Poisson equation Eqn. (14) and the definition of ${\bm{\Sigma}}_{\mu}({\bm{F}})$ (Eqn. (3)), we have
$\mu\left({{\bm{H}}}\right)=$ $\mu\left({{\mathrm{Q}}{\bm{G}}^{2}-({\mathrm{Q}}{\bm{G}})^{2}}\right)$
$=$ $\mu\left({{\bm{G}}^{2}-\left({{\bm{G}}-{\bm{F}}}\right)^{2}}\right)$
$=$ $\mu\left({{{\bm{F}}}{\bm{G}}+{\bm{G}}{\bm{F}}-{{\bm{F}}}^{2}}\right)$
$=$ $\mu\left({{{\bm{F}}}^{2}}\right)+\sum_{k=1}^{\infty}\mu\left({{{\bm{F}}}{\mathrm{Q}}^{k}{{\bm{F}}}+\left({{\mathrm{Q}}^{k}{{\bm{F}}}}\right){{\bm{F}}}}\right)$
$=$ ${\bm{\Sigma}}_{\mu}({\bm{F}}).$
@@
$\Big\|\left\|{\left\langle{{\bm{M}}}\right\rangle_{n}}\right\|^{1/2}\Big\|_{\mu,p}=$ $\sqrt{\left\|{\left\|{\sum_{k=1}^{n}{\bm{H}}(Z_{k-1})}\right\|}\right\|_{\mu,p/2}}$ (17)
$\leq$ $\sqrt{\left\|{n\left\|{\mu({\bm{H}})}\right\|+\left\|{\sum_{k=1}^{n}\left({{\bm{H}}(Z_{k-1})-\mu({\bm{H}})}\right)}\right\|}\right\|_{\mu,p/2}}$
$\leq$ $\sqrt{n\left\|{{\bm{\Sigma}}_{\mu}({\bm{F}})}\right\|+\left\|{\left\|{\sum_{k=0}^{n-1}\bar{\bm{H}}(Z_{k})}\right\|}\right\|_{\mu,p/2}}$
$\leq$ $\sqrt{n}\left\|{{\bm{\Sigma}}_{\mu}({\bm{F}})}\right\|^{1/2}+\left\|{\left\|{\sum_{k=0}^{n-1}\bar{\bm{H}}(Z_{k})}\right\|}\right\|_{\mu,p/2}^{1/2},$
where we denoted $\bar{{\bm{H}}}={{\bm{H}}}-\mu({\bm{H}})$.
@@
$\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\mu,p}\leq$ $C_{1}\sqrt{n(p\vee\log d)}\left\|{{\bm{\Sigma}}_{\mu}({\bm{F}})}\right\|^{1/2}+C_{1}\sqrt{p\vee\log d}\left\|{\left\|{\sum_{k=0}^{n-1}\bar{\bm{H}}(Z_{k})}\right\|}\right\|_{\mu,p/2}^{1/2}$ (18)
$+C_{2}(p\vee\log d)\left\|{\sup_{k\leq n}\|{\bm{X}}_{k}\|}\right\|_{\mu,p}+2\left\|{\left\|{{\bm{G}}(Z_{0})}\right\|}\right\|_{\mu,p},$
@@
If we assume Assumption 2.1 and ${\bm{F}}^{2}(z)\preccurlyeq{\bm{\Upsilon}}$, then by Lemma 4.2 (a), it holds that $\left({{\mathrm{Q}}{\bm{G}}^{2}}\right)(Z_{k-1})\preccurlyeq(64/9)t_{\operatorname{mix}}^{2}{\bm{\Upsilon}}$, thus $\langle{{\bm{M}}}\rangle_{n}\preccurlyeq(64/9)nt_{\operatorname{mix}}^{2}{\bm{\Upsilon}}$. By Theorem 2.1 (b), the inequality (16) still holds if we replace $d$ by $d_{n,p}=e\cdot r([(64/9)t_{\operatorname{mix}}^{2}{\bm{\Upsilon}}]\wedge[(1/n)\|\|\left\langle{{\bm{M}}}\right\rangle_{n}\|\|_{\mu,\frac{p}{2}}])$, and double the right-hand side. As $n\to\infty$, by ergodicity and the fact that $\mu({\bm{H}})={\bm{\Sigma}}_{\mu}({\bm{F}})$, we find that $d_{n,p}\to e\cdot r([(64/9)t_{\operatorname{mix}}^{2}{\bm{\Upsilon}}]\wedge\|{\bm{\Sigma}}_{\mu}({\bm{F}})\|)$, which is a natural dimension factor.
@@
    $\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\mu,p}\leq$ $\frac{16}{3}C_{1}\sqrt{(p\vee\log d)nt_{\operatorname{mix}}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}+\left({\frac{16}{3}C_{2}(p\vee\log d)+\frac{19}{3}}\right)t_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty},$
    Furthermore, if there exists an ${\bm{\Upsilon}}\in{\mathbb{H}}^{d}_{+}$ such that ${\bm{F}}^{2}(z)\preccurlyeq{\bm{\Upsilon}}$ for all $z\in{\mathcal{Z}}$, then the above inequality also holds when we replace $d$ by $e\cdot r([(64/9){\bm{\Upsilon}}]\wedge[(1/\lfloor{n}/{t_{\operatorname{mix}}}\rfloor)\|\|\langle{{\bm{M}}^{(t_{\operatorname{mix}})}}\rangle_{\lfloor{n}/{t_{\operatorname{mix}}}\rfloor}\|\|_{\mu,\frac{p}{2}}])$, and double the right-hand side, where ${\bm{M}}^{(t_{\operatorname{mix}})}$ is a matrix martingale defined in Appendix B.2.
@@
    $\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\mu,p}\leq$ $\frac{8}{3}C_{1}p\sqrt{(p\vee\log d)nt_{\operatorname{mix}}}\left[{\varkappa\mu(V)}\right]^{1/p}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{1/p}}$
    $+\left({\frac{16}{3}C_{2}(p\vee\log d)+\frac{11}{3}}\right)pt_{\operatorname{mix}}n^{1/p}\left[{\varkappa\mu(V)}\right]^{1/p}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{1/p}}.$
@@
$C_{1}\sqrt{p\vee\log d}\left\|{\left\|{\sum_{k=0}^{n-1}\bar{\bm{H}}(Z_{k})}\right\|}\right\|_{\mu,p/2}^{1/2}\lesssim\left({p\vee\log d}\right)^{3/4}t_{\operatorname{mix}}^{5/4}n^{1/4}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}+\left({p\vee\log d}\right)t_{\operatorname{mix}}^{3/2}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}.$ (19)
And it is easy to see that the other two residual terms are dominated by the term. If we further assume ${\bm{F}}^{2}(z)\preccurlyeq{\bm{\Upsilon}}$, then by Lemma 4.2 (a), the above inequality still holds if we replace $d$ introduced in the l.h.s. by $e\cdot r([(64/9)t_{\operatorname{mix}}^{2}{\bm{\Upsilon}}]\wedge[(1/n)\|\|\left\langle{{\bm{M}}}\right\rangle_{n}\|\|_{\mu,\frac{p}{2}}])$, and $d$ introduced in the r.h.s. (see Lemma 4.3) by $e\cdot r([(64/9)t_{\operatorname{mix}}^{2}{\bm{\Upsilon}}]\wedge[(1/\lfloor{n}/{t_{\operatorname{mix}}}\rfloor)\|\|\langle{{\bm{M}}^{(t_{\operatorname{mix}})}_{{\bm{H}}}}\rangle_{\lfloor{n}/{t_{\operatorname{mix}}}\rfloor}\|\|_{\mu,\frac{p}{4}}])$, and double the r.h.s., where the matrix martingale ${\bm{M}}^{(t_{\operatorname{mix}})}_{\bar{\bm{H}}}$ is defined in Appendix B.2 with $\bm{F}$} replaced by $\bar{\bm{H}}$.
@@
$\leq$ $\sup_{z\in{\mathcal{Z}}}\frac{[p\varkappa^{1/p}2^{3-1/p}/3](\mu(V)+V(z))^{1/p}t_{\operatorname{mix}}\|\|{\bm{F}}\|\|_{V^{1/p}}}{V^{1/p}(z)}$
$\lesssim$ $p\left[{\varkappa\mu(V)}\right]^{1/p}t_{\operatorname{mix}}\|\|{\bm{F}}\|\|_{V^{1/p}},$
@@
$C_{1}\sqrt{p\vee\log d}\left\|{\left\|{\sum_{k=0}^{n-1}\bar{\bm{H}}(Z_{k})}\right\|}\right\|_{\mu,p/2}^{1/2}\lesssim$ $p^{3/2}\left({p\vee\log d}\right)^{3/4}t_{\operatorname{mix}}^{5/4}n^{1/4}\left[{\varkappa\mu(V)}\right]^{2/p}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{1/p}}$
$+p^{3/2}\left({p\vee\log d}\right)t_{\operatorname{mix}}^{3/2}n^{1/p}\left[{\varkappa\mu(V)}\right]^{2/p}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{1/p}}.$
@@
$\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\xi,p}\leq\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\mu,p}+\frac{2}{n}\left({1+128(t_{\operatorname{mix}}/\log 4)^{p}(p+1)^{p+1/2}e^{-p}}\right)^{1/p}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{\infty}.$
@@
$\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\xi,p}\leq\frac{1}{n}\left\|{\sup_{k\leq n}\left\|{{\bm{S}}_{k}}\right\|}\right\|_{\mu,p}+\frac{32}{15n}\varkappa^{1/p}\left({\xi(V)+\mu(V)}\right)^{1/p}pt_{\operatorname{mix}}\left\|{\left\|{{\bm{F}}}\right\|}\right\|_{V^{1/p}}.$